OpenStudy (anonymous):

Simplify the following expression. 2x − 6y + 2z − 5x − 3y + 8z −3x − 9y + 10z −3x + 9y + 10z −3x + 9y − 10z 3x − 9y + 10z

OpenStudy (anonymous):

combine like terms: 2x-5x=-3x -6y-3y=-9y 2z+8z=10z So, -3x-9y+10z